# Quantum Kernel Methods under Scrutiny: A Benchmarking Study

### v2

@@ Acknowledgements.
This work was supported by the German Federal Ministry of Economic Affairs and Climate Action through the project AQUAS (grant no. 01MQ22003D), as well as the German Federal Ministry of Education and Research through the project H2Giga Degrad-EL3 (grant no. 03HY110D). Furthermore, the authors acknowledge Fraunhofer IAO for providing access to the GPU simulation cluster (CLARSA), which was established as part of the Competence Center Quantum Computing Baden-Württemberg. Here, J.S. is grateful to Vamshi Mohan Katukuri for his support with and excellent administration of the computing clusters. Furthermore, we would like to thank Moritz Link for his excellent contributions in an early stage of this work. Moreover, we are particularly grateful for insightful discussions with Roberto Flórez Ablan.
Appendices
Appendix A Some Background on Statistical Methods
@@
Appendix B Detailed Overview of Encoding Circuits
In the following we provide plots of each encoding circuits used in this study and briefly describe (cf. sQUlearn documentation) the corresponding features. Features are denoted with a feature vector x[i], while variationally trainable parameters ($\bm{\theta}$) are labeled p[i].
B.1 YZ_CX_EncodingCircuit
@@ B.9 ZZFeatureMap
Appendix C Note on Reproducibility of Results
@@
Appendix D Detailed Results
Here we support and complement the key findings of this study as presented in Sec. IV of the main text.
Figure A10: Analogous illustration to Fig. 4 with associated training performance scores. Results for each QKM and dataset are aggregated across data encoding circuits with corresponding optimal $n_{\mathrm{layers}}^{*}$, yielding minimum/maximum test performance scores for regression/classification, respectively. The upper panel displays corresponding findings for regression datasets. Here, (a) shows the Friedman dataset family with datsets for #$\#\mathrm{features}d=[5,15]$. In (b) we give results corresponding to the QFMNIST dataset family with datasets having number of principal components $d=[2,15]$. The lower panel shows the classification tasks. The results in (c) correspond to the two curves diff family with $d=4$ and degree $D=[2,20]$. In (d) we present the hidden manifold diff family with $d=4$ and manifold dimension $m=[2,20]$.
D.1 Details on Model Performance
In Sec. IV.1 we mention in connection with outliers in Fig. 4 that both regression and classification datasets exhibit a propensity for overfitting. The Friedman and QFMNIST regression datasets in Figs. A10 (a) and (b), respectively, reveal that each dataset shows lower whiskers with $\mathrm{MSE}\to 0$. Comparing to the corresponding test performance scores in Figs. 4 (a) and (b), this may also indicate signatures of overfitting especially for datasets with larger $d$. The corresponding training performance scores for the two curves diff and hidden manifold diff datasets are displayed in Figs. A10 (c) and (d), respectively. As discussed in the main text, since we investigate a large variety of combinations, it is unavoidable that some of them result in models that are too expressive for some tasks. Moreover, we point out that one should carefully investigate different training- and test set sizes in future studies.
D.2 Additions to Hyperparameter Correlation Analysis
To complement the report of correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameters themselves in Sec. IV.2, we provide a summary of all other dataset families not shown in the main text in Fig. A11.
@@ D.2 Additions to Hyperparameter Correlation Analysis
D.3 Details on the Influence of Encoding Circuits
When investigating the impact of data encoding circuits on QKM model performance in Sec. IV.3, we merely presented results of the Friedman (regression) and the two curves diff (classification) datasets in Fig. 8. Here, we provide analogue findings in Fig. A12 for the QFMNIST and the hidden manifold diff datasets.
@@ D.3 Details on the Influence of Encoding Circuits
D.4 Details on the Influence of Data Encoding Strategies
The results for testing the two different encoding strategies for the Friedman dataset for $d=5$ features as outlined in the main text, cf. Sec. IV.4 and Fig. 9, are displayed in Fig. A13. Contrary to the QFMNIST results in Fig. 10, in all cases, except for FQK, option 2, most accurate results are obtained for $n_{\mathrm{qubits}}=12$. Here, FQK and PQK approaches show similar behavior. The second encoding option for FQK appears to be less sensitive for the respective feature redundancy. Notably, that the redundancy does not have to be an integer multiple of the original number of features.
Figure A13: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK and PQK approaches for the Friedman regression dataset $d=5$ features. The results were aggregated over QKRR and QSVR approaches as well as over the corresponding encoding circuits, and then the median of the MSE test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
@@ D.4 Details on the Influence of Data Encoding Strategies
Figure A15 provides insights into the effect of the two different encoding strategies for the hidden manifold diff dataset with $m=13$. The findings discussed for the two curves diff dataset in Sec. IV.4 still apply.
Figure A15: Investigation of different encoding strategies as illustrated in Fig. 9 for FQK- and PQK-QSVC approaches, respectively for the hidden manifold diff classification dataset with manifold dimension $m=13$. Note, the feature dimension is kept constant at $d=4$ The results were aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
D.5 Details on the Analysis of PQK Design Options
@@
D.5.1 Overview on Test Performances
In the main text we investigate (semi-) partial correlations in Tabs. 1 and 2 between the choice of the outer kernel function $\kappa$ and corresponding hyperparameters as well as the measurement operator $O^{k}$ and resulting model test performances. In the following, we give a detailed insights into the test scores for each outer kernel (Gaussian, Matérn, RationalQuadratic) as a function of the measurement operator in Figs. A16 and A17 for the regression and classification tasks, respectively considered in this study. The data for each $O^{k}$ are aggregated across the encoding circuits considered here. Comparing these plots with the numbers listed in Tab. 1 helps to understand the reported statistically significant correlation. For instance, Fig. A16 (b) indicates that for each $O^{k}$ the Gaussian outer kernel leads to weaker test performance scores compared to the rest. This in turn reflects in the positive Spearman correlation coefficient in Tab. 1. Analogous conclusions can be drawn for the observed trends with respect to the different measurement operators considered in this work.
@@ D.5.1 Overview on Test Performances
Figure A17: ROC-AUC test score for QSVC-PQK models as a function of the measurement operators given in Eqs. (14),  (16) and (18) for Gaussian, Matér and RationalQuadratic outer kernel functions $\kappa$. This supports the discussion around Tab. 1 in Sec. IV.5 of the main text. The results corresponding to the following classification datasets are shown: (a) two curves diff dataset for $d=4$ and degree $D=13$. (b) hidden manifold diff dataset for $d=4$ and manifold dimension $m=13$. The data for each $O^{k}$ are aggregated across the encoding circuits considered in this work.
@@
D.5.2 Extended Correlation Analysis
In order to support the efforts in Sec. IV.5 to gain a deeper understanding of what are basic mechanisms in PQKs responsible for learning, we provide all (statistically significant) semi-partial correlations of the various hyerparameters of the QSVR/QSVC-PQK model for a given outer kernel $\kappa$, both among themselves and with model test performance in Tab. 2. Here, we note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.
Table 2: Statistically significant semi-partial correlations of the various hyperparameter of the QSVR/QSVC-PQK model for a given outer kernel function $\kappa$, both among themnselves and with test performance scores. This table supports the corresponding discussion in Ref. IV.5. We note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.

### v3

@@ 6 Conclusion
Acknowledgments This work was supported by the German Federal Ministry of Economic Affairs and Climate Action through the project AQUAS (grant no. 01MQ22003D), as well as the German Federal Ministry of Education and Research through the project H2Giga Degrad-EL3 (grant no. 03HY110D). Furthermore, the authors acknowledge Fraunhofer IAO for providing access to the GPU simulation cluster (CLARSA), which was established as part of the Competence Center Quantum Computing Baden-Württemberg. Here, J.S. is grateful to Vamshi Mohan Katukuri for his support with and excellent administration of the computing clusters. Furthermore, we would like to thank Moritz Link for his excellent contributions in an early stage of this work. Moreover, we are particularly grateful for insightful discussions with Roberto Flórez Ablan.
[…]
Availability of data and material The code developed for this study is available via GitLab [46]. Data used in this study is publicly available and can be accessed via the respective references provided in the manuscript. A guidance how to use the code and potentially reproduce the results is provided in the repository.
Declarations
[…]
Conflict of interest The authors declare that they have no conflict of interest.
Appendix A Detailed Theoretical Background
A.1 Details on Conventional Kernel Theory
The key idea behind the conventional kernelized approach to (supervised) machine learning is to find and analyze patterns by transforming the respective learning problem from the original input data domain $\mathcal{X}$ to a higher-dimensional (potentially infinite-dimensional) feature space $\mathcal{F}$, where the learning tasks can often be expressed in a linear form. This mapping is accomplished by a feature map $\phi:\mathcal{X}\to\mathcal{F};\mathbf{x}\mapsto\phi(\mathbf{x})$. Kernels, are real- or complex-valued symmetric and positive semi-definite functions of two input data points, $k:\mathcal{X}\times\mathcal{X}\to\mathbb{C}$. In this regard, another central concept is that of reproducing kernel Hilbert space (RKHS) which uniquely determines the kernel and vice versa [4] and in addition, for every kernel there exists at least one feature map such that [49]
Instead of explicitly computing the transformation $\phi(x)$ to the high-dimensional feature space, the kernel trick accomplishes the same result in the original input space through the kernel function. Besides, the Gaussian (RBF) kernel, we employ further common kernel functions in this work (specifically, for $\kappa$ in Eq. (7), i.e. the PQK definition). First, the Matérn kernel [44] with $\nu=\nicefrac{{3}}{{2}}$, i.e.
$\kappa^{\text{Mat}}_{3/2}(\mathbf{x},\mathbf{x}^{\prime})$ $=\left(1+\frac{\sqrt{3}\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert}{\ell}\right)$
×$\times\exp\left(-\frac{\sqrt{3}\lVert\mathbf{x}-\mathbf{x}^{\prime}\rVert}{\ell}\right)\,,$ (18)
and the RationalQuadratic kernel [15], i.e.,
$h(\mathbf{x})$ $=\sum_{i=1}^{N}c_{i}k(\mathbf{x},\mathbf{x}_{i})\,.$ (20)
Note that the number of terms in the sum is independent of the dimension of $\mathcal{H}_{k}$. The determination of the coefficients $\{c_{i}\}$ is a convex optimization problem [13]. This result is central to kernel methods used for supervised learning problems, e.g., SVMs [47], Gaussian Processes [44] or KRR [39]. All these algorithms are based on the kernel Gram matrix $K_{ij}=[k(\mathbf{x}_{i},\mathbf{x}_{j})]$, as detailed below for KRR and SVR.
A.2 KRR versus SVR
Kernel Ridge Regression (KRR) and Support Vector Regression (SVR) are both kernel-based regression techniques, but they differ in their formulation, optimization objectives, and practical applications. Given a dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$, KRR minimizes the following objective function [47]
$\min_{\mathbf{w}}\sum_{i=1}^{N}(y_{i}-\mathbf{w}^{T}\phi(x_{i}))^{2}+\lambda\lVert\mathbf{w}\rVert^{2}\,,$ (21)
where $\phi(x)$ is a feature map as introduced in Sec. 2, $\mathbf{w}$ is the weight vector in the transformed space, and $\lambda$ is a regularization parameter. Using the representer theorem, the solution takes the form Eq. (20), where the linear coefficient vecotr $\mathbf{c}$ is obtained by solving
$\mathbf{c}$ $=(\mathbf{K}+\lambda\mathbf{I})^{-1}\mathbf{y}\,.$ (22)
Therefore, the key characteristics of KRR are
  1. 1.
    Matrix inversion makes it computationally expensive, i.e. $\mathcal{O}(N^{3})$ for large datasets
  2. 2.
    The quadratic loss function penalizes large error more heavily than small ones
  3. 3.
    The Tikhonov regularization $\lambda$ balances smoothness and fit to data
The SVR aims to find a function $f(x)$ that predicts the targets $y$ with a margin of tolerance $\varepsilon$. For this, it uses the $\varepsilon$-intesive loss function and solves the following constrained optimization problem [47, 41]
$\min_{\mathbf{w},b,\xi,\xi^{*}}\frac{1}{2}\lVert\mathbf{w}\rVert^{2}+C\sum_{i=1}^{N}(\xi_{i}+\xi_{i}^{*})\,,$ (23)
subject to the constraints
$y_{i}-\mathbf{w}^{T}\phi(x_{i})-b$ $\leq\varepsilon+\xi_{i}\,,$ (24)
$\mathbf{w}^{T}\phi(x_{i})+b-y_{i}$ $\leq\varepsilon+\xi_{i}^{*}\,,$ (25)
$\xi_{i},\xi_{i}^{*}$ $\geq 0\,.$ (26)
Here, $\varepsilon$ is a threshold margin for error tolerance, $\xi_{i},\xi_{i}^{*}$ are slack variables that allow deviations beyond $\varepsilon$, and $C>0$ is the regularization parameter controlling the tradeoff between margin and error. The dual formulation leads to the prediction function
$f(x)$ $=\sum_{i=1}^{N}(\alpha_{i}-\alpha_{i}^{*})K(x,x_{i})+b\,,$ (27)
where $\alpha_{i},\alpha_{i}^{*}$ are the dual coefficients. The key characteristics of SVR are
  1. 1.
    The $\varepsilon$-intensive loss encourages sparsity and robustness to outliers
  2. 2.
    Solving the quadratic problem can be also computationally demanding, i.e. $\mathcal{O}(N^{2})$ to $\mathcal{O}(N^{3})$, but is often more scalable than KRR for large datasets
To this end, the choice between KRR and SVR depends on the dataset size, the desired model complexity, and the tradeoff between computational efficiency and model performance. KRR is well-suited for small datasets with Gaussian noise, while SVR employs a sparse optimization and is more robust to outliers as well as better suited for large datasets.
We refer to Refs. [47, 44, 39] for a thorough introduction into kernel theory and kernel methods.
Appendix B Some Background on Statistical Methods
@@
Appendix C Detailed Overview of Encoding Circuits
In the following we provide plots of each encoding circuits used in this study and briefly describe (cf. sQUlearn documentation) the corresponding features. Features are denoted with a feature vector x[i], while variationally trainable parameters ($\boldsymbol{\theta}$) are labeled p[i].
C.1 YZ_CX_EncodingCircuit
@@ C.9 ZZFeatureMap
Appendix D Details on Datasets
[…]
Friedman The Friedman #1 regression problem is described in Refs. [17, 9]. The dataset consists of $d\geq 5$ independent features distributed uniformly on the interval $[0,1]$. The labels $y$ satisfy
Appendix E Details on QKMTuner
As indicated in Sec. 3.3, we developed the tool QKMTuner to facilitate the extensive hyperparameter search of this study. Its core functionalities are sketched in Fig. 2 consisting of two main routines: a hyperparameter optimization within a grid-search and a hyperparameter importance analysis. The grid-search is given a list of data encoding circuits for each of which it builds a user-defined $n_{\mathrm{qubits}}\times n_{\mathrm{layers}}$-grid on which a hyperparameter search is performed for each grid point and each data encoding circuit for a given dataset. For evaluating hyperparameter importances, the respective method only takes a list of data encoding circuits and automatically searches for the optimal hyperparameters including $n_{\mathrm{qubits}}$ and $n_{\mathrm{layers}}$. In both cases, the hyperparameters are determined by maximizing the minimum between the mean and median of five-fold cross-validation scores. This choice of the objective function prevents the scores from becoming too optimistic and thus helps to prevent overfitting towards an easy fold. The respective scoring method can be set manually, whereas in this work we use the area under the receiver operating characteristic curve (ROC-AUC) from prediction scores for classification and the (negative) mean squared error (MSE) for regression tasks, respectively. For hyperparameter sampling we use the tree-structured parzen estimator [58], albeit different algorithms as provided by Optuna are supported.
Appendix F Note on Reproducibility of Results
@@
Appendix G Detailed Results
Here we support and complement the key findings of this study as presented in Sec. 4 of the main text.
Figure A10: Analogous illustration to Fig. 4 with associated training performance scores. Results for each QKM and dataset are aggregated across data encoding circuits with corresponding optimal $n_{\mathrm{layers}}^{*}$, yielding minimum/maximum test performance scores for regression/classification, respectively. For comparison, we provide classical KRR/SVR and SVC results each based on a RBF kernel. Corresponding hyperparameters are also optimized using Optuna with parameter ranges chosen as given in Sec. 3.3. The upper panel displays corresponding findings for regression datasets. Here, (a) shows the Friedman dataset family with datsets for #$\#\mathrm{features}d=[5,15]$. In (b) we give results corresponding to the QFMNIST dataset family with datasets having number of principal components $d=[2,15]$. The lower panel shows the classification tasks. The results in (c) correspond to the two curves diff family with $d=4$ and degree $D=[2,20]$. In (d) we present the hidden manifold diff family with $d=4$ and manifold dimension $m=[2,20]$.
G.1 Details on Model Performance
In Sec. 4.1 we mention in connection with outliers in Fig. 4 that both regression and classification datasets exhibit a propensity for overfitting. The Friedman and QFMNIST regression datasets in Figs. A10 (a) and (b), respectively, reveal that each dataset shows lower whiskers with $\mathrm{MSE}\to 0$. Comparing to the corresponding test performance scores in Figs. 4 (a) and (b), this may also indicate signatures of overfitting especially for datasets with larger $d$. The corresponding training performance scores for the two curves diff and hidden manifold diff datasets are displayed in Figs. A10 (c) and (d), respectively. As discussed in the main text, since we investigate a large variety of combinations, it is unavoidable that some of them result in models that are too expressive for some tasks. Moreover, we point out that one should carefully investigate different training- and test set sizes in future studies.
G.2 Additions to Hyperparameter Correlation Analysis
To complement the report of correlation analyses between QKM hyperparameters and test performance scores, as well as between the hyperparameters themselves in Sec. 4.2, we provide a summary of all other dataset families not shown in the main text in Fig. A11.
@@ G.2 Additions to Hyperparameter Correlation Analysis
G.3 Details on the Influence of Encoding Circuits
When investigating the impact of data encoding circuits on QKM model performance in Sec. 4.3, we merely presented results of the Friedman (regression) and the two curves diff (classification) datasets in Fig. 8. Here, we provide analogue findings in Fig. A12 for the QFMNIST and the hidden manifold diff datasets.
@@ G.3 Details on the Influence of Encoding Circuits
G.4 Influence of Data Encoding Strategies
As already implicitly indicated in Sec. 4.2, it is known that the number of qubits and layers of data encoding circuits determine the Fourier frequency spectrum of the resulting quantum model and that encoding features redundantly can help in getting more accurate results [49, 51]. However, from a practical point of view it is not clear how to best distribute the features of a data point among the available qubits in a particular encoding circuit. To shed some light on this aspect, we compare QKMs built from selected encoding circuits following the two encoding schemes illustrated in Fig. A13. In both embedding schemes, features are encoded in the qubits from top to bottom. Differences in the two encoding strategies occur when the number of qubits is larger than the number of features. In “option 1”, after all features have been encoded once in a layer, the encoding is repeated and then cut off once the final qubit is reached. In “option 2”, the enumeration of features is not reset at the next layer. As a consequence, each feature in “option 1” is always assigned to the same qubit, whereas in “option 2” the features can be shuffled across qubits.
Figure A13: Schematic illustration of two different data encoding options considered in the investigations of Figs. A14 and  A17.
Using the QFMNIST dataset with $d=5$ as an example, Fig. (A14) shows the findings for this regression tasks. The results for each quantum kernel are aggregated across the regression methods and the respective encoding circuits. Then, the median of the resulting MSE test score is calculated. For FQK, both embedding options show that largest scores are achieved for integer multiples of the number of features ($n_{\mathrm{qubits}}=10$). It can be seen that higher scores occur more frequently with embedding “option 2”. For PQK, it appears that no feature redundancies are necessary in either case. Here, however, embedding “option 1” often delivers larger scores.
Figure A14: Investigation of different encoding strategies as illustrated in Fig. A13 for FQK and PQK approaches for the QFMNIST regression dataset with $d=5$ principal components. The results are aggregated over QKRR and QSVR approaches as well as over the corresponding encoding circuits, and then the median of the MSE test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
The results for testing the two different encoding strategies for the Friedman dataset for $d=5$ features are displayed in Fig. A15. Contrary to the QFMNIST results in Fig. A14, in all cases, except for FQK, option 2, most accurate results are obtained for $n_{\mathrm{qubits}}=12$. Here, FQK and PQK approaches show similar behavior. The second encoding option for FQK appears to be less sensitive for the respective feature redundancy. Notably, that the redundancy does not have to be an integer multiple of the original number of features.
Figure A15: Investigation of different encoding strategies as illustrated in Fig. A13 for FQK and PQK approaches for the Friedman regression dataset $d=5$ features. The results were aggregated over QKRR and QSVR approaches as well as over the corresponding encoding circuits, and then the median of the MSE test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
@@ G.4 Influence of Data Encoding Strategies
Figure A17 displays the results for the two curves diff classification dataset for $D=13$ (with $d=4$). Again, results are aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores is evaluated. It can be seen that in contrast to the previous example, both embedding options deliver the best classification scores for non-integer multiples of the number of features (for PQK and FQK). Here, “option 1” performs better in each case. In addition, larger values of $n_{\text{layers}}$ and $n_{\text{qubits}}$ lead to increasing classification scores. The difference to the previous example (Fig. A14) might be due to significantly larger dataset complexity. This also explains the overall larger difference in performance scores for the two curves diff.
Figure A17: Investigation of different encoding strategies as illustrated in Fig. A13 for FQK- and PQK-QSVC approaches, respectively for the two curves diff classification dataset with $D=13$. Note, the feature dimension is kept constant at $d=4$. The results are aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores is calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
Figure A18 provides insights into the effect of the two different encoding strategies for the hidden manifold diff dataset with $m=13$. The findings discussed previously for the two curves diff dataset still apply.
Figure A18: Investigation of different encoding strategies as illustrated in Fig. A13 for FQK- and PQK-QSVC approaches, respectively for the hidden manifold diff classification dataset with manifold dimension $m=13$. Note, the feature dimension is kept constant at $d=4$ The results were aggregated over the respective encoding circuits and then the median of the ROC-AUC test scores was calculated in each case. The upper panel corresponds to FQKs, while the lower panel shows PQKs, whereas the left column represents “option 1” and the right column “option 2”, respectively. As such, we have: (a) FQK, option 1, (b) FQK, option 2, (c) PQK, option 1 and (d) PQK, option 2.
G.5 Details on the Analysis of PQK Design Options
@@
G.5.1 Overview on Test Performances
In the main text we investigate (semi-) partial correlations in Tabs. 1 and 2 between the choice of the outer kernel function $\kappa$ and corresponding hyperparameters as well as the measurement operator $O_{k}$ and resulting model test performances. In the following, we give a detailed insights into the test scores for each outer kernel (Gaussian, Matérn, RationalQuadratic) as a function of the measurement operator in Figs. A19 and A20 for the regression and classification tasks, respectively considered in this study. The data for each $O_{k}$ are aggregated across the encoding circuits considered here. Comparing these plots with the numbers listed in Tab. 1 helps to understand the reported statistically significant correlation. For instance, Fig. A19 (b) indicates that for each $O_{k}$ the Gaussian outer kernel leads to weaker test performance scores compared to the rest. This in turn reflects in the positive Spearman correlation coefficient in Tab. 1. Analogous conclusions can be drawn for the observed trends with respect to the different measurement operators considered in this work.
@@ G.5.1 Overview on Test Performances
Figure A20: ROC-AUC test score for QSVC-PQK models as a function of the measurement operators given in Eqs. (10),  (12) and (14) for Gaussian, Matér and RationalQuadratic outer kernel functions $\kappa$. This supports the discussion around Tab. 1 in Sec. 4.4 of the main text. The results corresponding to the following classification datasets are shown: (a) two curves diff dataset for $d=4$ and degree $D=13$. (b) hidden manifold diff dataset for $d=4$ and manifold dimension $m=13$. The data for each $O_{k}$ are aggregated across the encoding circuits considered in this work.
@@
G.5.2 Extended Correlation Analysis
In order to support the efforts in Sec. 4.4 to gain a deeper understanding of what are basic mechanisms in PQKs responsible for learning, we provide all (statistically significant) semi-partial correlations of the various hyerparameters of the QSVR/QSVC-PQK model for a given outer kernel $\kappa$, both among themselves and with model test performance in Tab. 2. Here, we note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.
Table 2: Statistically significant semi-partial correlations of the various hyperparameter of the QSVR/QSVC-PQK model for a given outer kernel function $\kappa$, both among themnselves and with test performance scores. This table supports the corresponding discussion in Ref. 4.4. We note that while negative correlation coefficients w.r.t. the MSE point at overall better model performance, it is the other way round for correlations with ROC-AUC scores.